Here's another one for you.I promise this will be it for today...HA HA!!
Erin's age is 3 times Warren's age. In 4 years she will be twice as old as he will be. How old is each now?
Thanks for your help.
Joanne

Answer by scott8148(6628)

E=3W __ E+4=2(W+4)

substituting gives (3W)+4=2(W+4) __ 3W+4=2W+8 __ W=4